TPC status

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Sun Jan 28 2001 - 16:07:23 EST

  • Next message: Peter H. L. Christiansen: "BRAT-1-11-4"

     
    Dear Collaborator,
    
    In preparation for the upcoming long run several hardware tasks has to be completed. The highest priority is the repairs of the TPC with which this message is concerned. I plan shortly to give you information on the other action items, notably the Drift Chambers, small beam-beam counters, but also installation of additional TOFW panels, calibration chambers and repair of C1.
    
    At the collaboration meeting, and in subsequent discussions by e-mail the action for modifications of the TPC was laid out and agreed upon. Since before Christmas,  Ramiro, Ed Baker, and Chellis has been working on this important project. Additionally Jens Ivar who will be resident at BNL for 10 month will be part of this effort in testing with cosmic rays either in the IR or in the lab.
    
     Once the repairs have been complete and the cosmic rays testing done, the TPCs will be re-installed in the spectrometer arm, and checkout and tuning with gating grid pulser will be done.
    
    I will thank Ramiro for supplying the report enclosed in this message
    
    Best regards
    
                Flemming
    
     
    
    Status Report
    
    TPM1: 
                All pieces necessary for winding new anode plane are ready. 
                New HV bus is being put together. 
                Machining to remove the old HV bus will proceed next. 
                Winding follows. 
    
               Once wire plane is ready we will setup a cosmic ray trigger. (this could be done at the IR or at the lab. depending on convenience) 
    
    TPM2: 
                Big gas leak was identified where entrance-exit foils became separated from double sided tape. This failure of the tape is common to all new TPCs. 
                To solve that problem we will remove the old foils and the tape and replace them with a system of frames and an O-ring. 
    
                Additional leaks were found in some of the plated thru holes that connect the pads to the connectors. This will be fixed with epoxy. 
    
    T1: 
               Leaks were found where foils come apart from double sided tape. Big leaks were also present at the junction of bottom plate and rest of lucite box. O-ring groove looks suspicious. 
               We also decided to replace the windows. Space limitations do not allow for similar solutions as for TPM2. The solution will be to glue the foil and keep it in place with a thin bolted frame. 
               The TPC was moved to the lab and work is in progress to replace first resistor in resistor chain. The value of all resistors in the T1 resistor chain is 5.2 M ohms.
     
    
    T2: 
             The TPC will be removed from platform and leak tested with Ar isobutane mixture. Windows will be replaced and, if necessary, first resistor changed. There is a crack in one of the lucite pieces, we have to see if it may grow. 
    
     
    
     
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    



    This archive was generated by hypermail 2b29 : Sun Jan 28 2001 - 16:01:54 EST